Given the matrices (Each matrices are inside a [ ] )
A= 3 0 , B= 4 -1 , C= 1 4 2 , D= 1 5 2 , E= 0 1 3
-1 2 0 2 3 1 5 -1 0 1 -1 1 2
1 1 3 2 4 4 1 3

1.-4C^2
2.(E-D)^T
3. A(BC)
4. (B^T B)C
5. 3B-AB
1
Expert's answer
2020-06-02T17:08:09-0400


A= "\begin{bmatrix}
 3 & 0 \\
 -1 & 2 \\
 1 & 1
\end{bmatrix}" , B = "\begin{bmatrix}
 4 & -1 \\
 0 & 2
\end{bmatrix}"

C = "\begin{bmatrix}
 1 & 4 & 2 \\
 3 & 1 & 5 \\
\end{bmatrix}" , D = "\begin{bmatrix}
 1 & 5 & 2 \\
 -1 & 0 & 1 \\
 3 & 2 & 4
\end{bmatrix}" ,

E= "\begin{bmatrix}
 0 & 1 & 3 \\
 -1 & 1 & 2 \\
 4 & 1 & 3
\end{bmatrix}"

1. Since C is not a square matrix , C² doesn't exist. Hence -4C² doesn't exist

2. E - D = "\begin{bmatrix}
 -1 & -4 & 1 \\
 0 & 1 & 1 \\
 1 & -1 & -1
\end{bmatrix}"

So (E-D)^T = "\begin{bmatrix}
 -1& 0 & 1 \\
 -4 & 1 & -1 \\
 1 & 1 & -1
\end{bmatrix}"

3. BC = "\begin{bmatrix}
 4 & -1 \\
 0 & 2
\end{bmatrix}" X "\begin{bmatrix}
 1 & 4 & 2 \\
 3 & 1 & 5 \\
\end{bmatrix}"

= "\begin{bmatrix}
 1 & 15 & 3 \\
 6 & 2 & 10 \\
\end{bmatrix}"


4. B^T = "\begin{bmatrix}
 4 & 0 \\
 -1 & 2
\end{bmatrix}"

B^TB = "\begin{bmatrix}
 4 & 0 \\
 -1 & 2
\end{bmatrix}" "\begin{bmatrix}
 4 & -1 \\
 0 & 2
\end{bmatrix}"

= "\begin{bmatrix}
 16 & -4 \\
 -4 & 5
\end{bmatrix}"

(B^TB)C = "\begin{bmatrix}
 16 & -4 \\
 -4 & 5
\end{bmatrix}" "\begin{bmatrix}
 1 & 4 & 2 \\
 3 & 1 & 5 \\
\end{bmatrix}"

= "\begin{bmatrix}
 4 & 60 & 12 \\
 11 & -11 & 17 \\
\end{bmatrix}"

5. 3B is of order 2x2 and AB is of order 3x2

So 3B - AB is not conformable for subtraction